Can you help and show me if I have done this problem correctly. Williamston Daily a daily newspaper. When they first started printing three years ago they had 6000 home delivery subscribers. The number of subscribers increases by 69 every month. Write an equation that shows the relationship between the number of subscribers, B, and the number of months the paper has been printed, n.
I think you would write this like, B = 6000n + 69 am I correct or not.
Thanks :o)

Answer by Theo(13342)

3 years ago they have 6000 subscribers
the number of subscribers increases by 69 each month.

let n = number of months the paper has been printed.

formula would be:

B = 6000 + 69*n

when n = 0, B = 6000
when n = 1, B = 6000 + 69*1`
when n = 2, B = 6000 + 69*2
etc.

after 3 years, the number of subscribers becomes:

6000 + 69*36 = 6000 + 2484 which equals 8484.